4) Tolerances for a new part require weights between 248.1 and 268.5 pounds. The part is made using a process that has a mean of 257.5 pounds with a standard deviation of 2.3 pounds. The process population is normally distributed.

What is the sigma capability for this process?

(use four decimal places)

5)

Tolerances for a new part require weights between 241.8 and 266.4 pounds. The part is made using a process that has a mean of 253.1 pounds with a standard deviation of 2.6 pounds. The process population is normally distributed.

If the process can be adjusted so that it is centered, what will be the sigma capability for the adjusted process?

(use four decimal places)

6)

Requirements for the width of a tractor engine component are 23.85 +/- 0.603 millimeters. The current process produces components with an average width of 23.488 and a population standard deviation of 0.094. The process is normally distributed.

What is the sigma capability of this process?

(Use four decimal places)

8)

Requirements for the width of a tractor engine component are 23.46 +/- 0.636 millimeters. The current process produces components with an average width of 23.534 and a population standard deviation of 0.085. The process is normally distributed.

If the control limits are set at +/- 2 standard deviations instead of +/- 3 standard deviations from the mean, what is the Cpk of this process?

(Use four decimal places)
Answer #1

Answering the first question as HOMEWORKLIB policy

Question 4: Tolerances for a new part require weights between 248.1 and 268.5 pounds. The part is made using a process that has a mean of 257.5 pounds with a standard deviation of 2.3 pounds. The process population is normally distributed. What is the sigma capability for this process? (use four decimal places)

Answer: Lower Specification = 248.10, Upper Specification = 268.50, Mean = 257.50, SD = 2.30

Since the data is normally distributed,

Sigma Level = (Upper specification - Mean) / SD = (268.50 - 257.50) / 2.30 = 4.7826

Also,

Sigma Level = (Mean - Upper specification) / SD = (257.50 - 248.10) / 2.3 = 4.0870

Either of the two (4.7826 or 4.0870) can be considered as the Sigma Level.